OpenStudy (anonymous):

Choose the correct right-hand side to make the equation an identity. cos(x)/1+csc(x)-cos(x)/1-csc(x)=

zepdrix (zepdrix):

Kind of hard to read with the way it's formatted. Is this what we're dealing with? :)\[\large\rm \frac{\cos x}{1+\csc x}-\frac{\cos x}{1-\csc x}\]

zepdrix (zepdrix):

`Choose the correct right-hand side` Were you given some choices?